# The AI Memory Gap: Users Misremember What They Created With AI or Without

### v1

@@ 4.4.1. Idea component
For ideas, the model estimated a hit rate of $d_{\text{idea}}=.86$ (95% HDI $[.84,.88]$), indicating that the vast majority of studied ideas were correctly recognised as old. Conditional on a hit, the true source was recalled on $s_{\text{idea}}=.39$ of trials ($[.36,.43]$), so source memory accompanied about two out of five detections. When the source was forgotten, participants guessed “AI” with probability $\beta_{\text{idea}}=.43$ ($[.39,.47]$), indicating a bias toward guessing noAI (the HDI excludes .50.50). The false-alarm rate for lure ideas was $f_{\text{idea}}=.29$ ($[.26,.33]$), reflecting a somewhat liberal old–new criterion.
@@ 4.4.2. Elaboration component
Elaborations yielded the same high hit rate as ideas, $d_{\text{elab}}=.86$ ($[.84,.88]$), but with a markedly higher source-memory probability, $s_{\text{elab}}=.62$ ($[.57,.65]$), showing that source was correctly identified on well over half of detections. When source memory failed, participants showed a bias toward the AI: $\beta_{\text{elab}}=.60$ ($[.56,.64]$), with the HDI clearly above .50.50. The false-alarm rate was again $f_{\text{elab}}=.29$ ($[.26,.33]$), closely matching the idea component.
@@ 4.4.4. Summary
Taken together, the results reveal a clear memory advantage for elaboration sources ($s_{\text{elab}}>s_{\text{idea}}$) and a selective AI guessing bias on elaborations ($\beta_{\text{elab}}>.50$ and $\beta_{\text{idea}}$), while overall recognition sensitivity and response criterion are effectively identical for ideas and elaborations. Accordingly, we interpret AI-related differences as shifts in source monitoring and elevated $\beta_{\text{elab}}$, not as changes in item discriminability.
@@ 4.6. Summary
  • AI involvement reduced item memory accuracy, dropping from 97.6% without any AI to 87.9% for withAI/withAI. This corresponds to substantially lower odds of correct recognition with AI use – especially at the elaboration stage (94.7%, OR = 0.30), compared to a smaller reduction for AI-ideation (97.3%, OR = 0.61).
@@ 5. Discussion
Moreover, source attribution accuracy was lowest in mixed workflows (AI support for either ideation or elaboration), indicating that keeping workflows with AI consistent supports memory – although this still did not reach accuracies of not using AI (RQ2).
MPT modelling confirmed that the effects were not due to diminished sensitivity to old items. It confirmed weaker source memory for ideas than elaborations, and revealed different guessing biases when sources were forgotten—towards self for ideas and towards AI for elaborations.
Finally, confidence ratings mirrored these patterns: They were highest without any AI. People overall gave rather high absolute confidence ratings, even when being wrong, and slightly overestimated their overall performance at the end. This indicates that they tended to be overconfident about their own performance (RQ1, RQ3).
@@ 5.1.2. Design
These theoretical insights have clear implications for design. AI outputs often lack the perceptual and contextual cues that the SMF identifies as essential for accurate source attribution, making it especially easy to misremember their origin. To directly support memory in human-AI co-creative processes, interaction and UX design should embed salient cues that help users retain a “memory of self-generation.” For example, designs might explore using varying animations, colours, interaction modes, etc. during content generation, which are distinct between own creation vs AI use.
@@ 5.2. Principles of Human-AI Interaction in the Context of the AI Memory Gap
Our findings show that consistent workflows support more accurate memory. Systems where correct attribution might be as important as the final output should therefore support consistent interactions. Concretely, users should be free to involve AI at every stage of the creative process, but they should not be forced or implicitly nudged to rely on AI for any particular task or subtask. When source attribution matters, users may choose to complete each sequence from ideation through elaboration either entirely on their own or entirely with AI. Consistency could also be established by always delegating a specific stage to AI (e.g., always ideate together with AI), while completing the remaining stages alone. And finally, evaluators should recognise that irregular processes are more prone to memory errors.
@@
While some systems make AI contributions more transparent (e.g., Hoque et al. (2024)), contributions to human thought processes are far harder to capture and later recall. Designers might experiment with tools that record prompts and drafts together, creating a map of thoughts alongside the evolving ideas and making AI’s influence on cognitive processes more visible. However, evaluators should acknowledge that source attributions will always be incomplete and avoid demanding certainty where it cannot exist. Practitioners should be aware that AI may shape not only their writing but also their thinking, often in ways that are difficult to recognize, even for the author themselves.
@@ 5.3.2. Memory vs Text Cues
5.4. Limitations, Generalizability and Future work
This study comes with several limitations:
First, we only examined the impact of AI usage during ideation and elaboration tasks as part of the writing process. We opted for those tasks as they represent fundamental stages that precede and influence many other creative processes. Whether our findings apply to other domains, such as coding, image and video creation, or design, needs to be explored in future work. Related, our experimental task mimicked a co-creative workflow for all AI conditions, where the human always made the final decision. Situations in which AI makes executive decisions on its own remain open for investigation. Based on our results, we hypothesise that this would further reduce people’s ability to remember parts of the creative process.
Furthermore, our study focused on a small subset of possible AI interactions, specifically using an LLM (GPT-4o) through a chatbot interface. Using a different type of AI, a different model, or an alternative mode of interaction, such as inline suggestions, could lead to different outcomes.
In addition, people in our study created ideas in one session, whereas real-world creative workflows often stretch over multiple sessions. We also tested people’s source memory after a fixed time gap of approximately one week, and it remains unclear how shorter or longer retention intervals might influence memory performance in this context.
@@ A.1. Problem Statements
The known-topic and unknown-topic distractors can be found in our project repository (https://osf.io/anonymous/).

### v2

@@ 4.4.1. Idea component
For ideas, the model estimated a hit rate of $d_{\text{idea}}=.86$ (95% HDI $[.84,.88]$), indicating that the vast majority of studied ideas were correctly recognised as old (i.e. ideas the participant had worked on in Phase 1). Conditional on a hit, the true source was recalled on $s_{\text{idea}}=.39$ of trials ($[.36,.43]$), so source memory accompanied about two out of five detections. When the source was forgotten, participants guessed “AI” with probability $\beta_{\text{idea}}=.43$ ($[.39,.47]$), indicating a bias toward guessing noAI (the HDI excludes .50.50). The false-alarm rate for lure ideas was $f_{\text{idea}}=.29$ ($[.26,.33]$), reflecting a somewhat liberal old–new criterion.
@@ 4.4.2. Elaboration component
Elaborations yielded the same high hit rate as ideas, $d_{\text{elab}}=.86$ ($[.84,.88]$), but with a markedly higher source-memory probability, $s_{\text{elab}}=.62$ ($[.57,.65]$), showing that source was correctly identified on well over half of detections. When source memory failed, participants showed a bias toward AI: $\beta_{\text{elab}}=.60$ ($[.56,.64]$), with the HDI clearly above .50.50. The false-alarm rate was again $f_{\text{elab}}=.29$ ($[.26,.33]$), closely matching the idea component.
@@ 4.4.4. Summary
Taken together, the results reveal a clear memory advantage for elaboration sources ($s_{\text{elab}}>s_{\text{idea}}$) and a selective AI guessing bias on elaborations ($\beta_{\text{elab}}>.50$ and $\beta_{\text{idea}}$), while overall recognition sensitivity and response criterion are effectively identical for ideas and elaborations. Accordingly, we interpret AI-related differences as reflecting changes in source monitoring and a stronger bias to attribute elaborations to AI ($\beta_{\text{elab}}$), not as changes in item recognition or discriminability (i.e., between old and new items, see Section 4.4).
@@ 4.6. Summary
  • AI involvement reduced item memory accuracy, dropping from 97.6% without AI to 87.9% for withAI/withAI. This corresponds to substantially lower odds of correct recognition with AI, especially at the elaboration stage (94.7%, OR = 0.30), compared to a smaller reduction for AI-ideation (97.3%, OR = 0.61).
@@ 5. Discussion
Moreover, source attribution accuracy was lowest in mixed workflows (AI support for either ideation or elaboration), indicating that keeping workflows with AI consistent supports memory – although this still did not reach the accuracy of not using AI (RQ2).
MPT modelling confirmed that the effects were not due to diminished sensitivity to old items. It confirmed weaker source memory for ideas than elaborations, likely because only the elaborations were shown verbatim, giving them richer, more distinctive, and perceptually detailed memory cues. This interpretation aligns with the SMF, which predicts that externally presented materials (such as the visible elaborations) provide more perceptual and contextual detail for memory to work with than internally retrieved content (such as the underlying ideas or thoughts), giving them a stronger basis for later source judgments. We contextualise this in Section 5.3.2.
Our MPT results further revealed different guessing biases when sources were forgotten – towards self for ideas and towards AI for elaborations. From an SMF perspective, this asymmetry might be explained by differences in the qualitative characteristics used during source monitoring. Ideas relate to internal cognitive operations and often lack distinctive perceptual detail; when memory is weak, participants may therefore rely on heuristic cues that make such abstract, thought-like content (i.e. underlying ideas) feel self-generated; an effect that may be amplified by having further developed these ideas during the elaboration task. In contrast, elaborations consist of fluent and sentence-level text-features that participants may associate with external sources, such as our AI chatbot, leading to an AI-biased guess when source memory fails.
Finally, confidence ratings mirrored these patterns: They were highest without any AI. People overall gave rather high absolute confidence ratings, even when being wrong, and slightly overestimated their overall performance at the end. This indicates that they tended to be overconfident about their own performance (RQ1, RQ3). At the same time, the decrease in confidence in AI-supported workflows likely reflects an implicit metacognitive sensitivity to weaker or less distinctive memory cues in these conditions. This suggests a limited awareness that source attribution is more difficult in human-AI workflows, even though confidence still exceeded actual accuracy and participants did not reliably identify when AI had been used.
@@ 5.1.2. Design
These theoretical insights have clear implications for design. AI outputs often lack the perceptual and contextual cues that the SMF identifies as essential for accurate source attribution, making it especially easy to misremember their origin. To directly support memory in human-AI co-creative processes, interaction and UX design should embed salient cues that help users retain a “memory of self-generation.” For example, designs might explore using varying animations, colours, interaction modes, etc. during content generation, which are distinct between own creation vs AI use. More broadly, future designs could explore leveraging the concepts of UI fragmentation (Buschek, 2024) and seamful design (Inman and Ribes, 2019) in context of the design dimensions for AI writing tools by Lee et al. (2024) to achieve such cues.
@@ 5.2. Principles of Human-AI Interaction in the Context of the AI Memory Gap
Our findings show that consistent workflows support more accurate memory. Users who rely on their memory, such as when systems provide limited provenance, benefit when a sequence from ideation through elaboration is completed either entirely on their own or entirely with AI. More consistency can also be achieved by setting “boundaries” and “limiting AI to specific tasks”, although this might only apply to a subgroup of writers who want to retain authenticity, a feeling of ownership and control, or enjoyment of the craft (Guo et al., 2025). In practice, this might be difficult to realise for many as AI is often integrated in flexible and opportunistic ways. Still, the principle can be supported by keeping certain aspects of writing workflows consistent: Even if users mix AI into different stages of their process, the system itself should stay consistent. Critically, current AI writing assistants (e.g., ChatGPT) tend to offer a “next step”. For example, when asked to proofread text, the AI may also suggest conceptual improvements, producing an output that mixes contributions across idea and elaboration stages. Keeping the AI’s role consistent, even as users employ it at different moments, may therefore support better memory performance. These stable roles could be reinforced through consistent interface designs, such as placing idea-generation features in a separated UI view, while editing assistance occurs in inline interactions. Finally, evaluators should recognise that inconsistent processes are more prone to memory errors.
@@
While some systems make AI contributions more transparent (e.g., Hoque et al. (2024)), contributions to human thought processes are far harder to capture and later recall. Designers might experiment with tools that record prompts and drafts together, creating a map of thoughts alongside the evolving ideas and making AI’s influence on cognitive processes more visible. However, evaluators should acknowledge that source attributions will always be incomplete and avoid demanding certainty where it cannot exist. Practitioners should be aware that AI may shape not only their writing but also their thinking, often in ways that are difficult to recognize, even for the authors themselves.
Contextualisation with other principles
We contextualise these principles with two widely recognised sets of human-AI interaction guidelines in the literature. First, the 18 guidelines by Amershi et al. (2019) are structured along a timeline from “initially” to “during interaction” to “over time”. While human memory is not explicitly considered, it might best fit with design choices in line with guideline 12, which says that systems should ”Remember recent interactions”. Moreover, our work motivates an additional category: “after active use”. Second, we relate our work to the six design principles for generative AI applications by Weisz et al. (2024). While these do not mention memory, our insights could extend their principles of “Design for Mental Models” or “Design for Imperfection”, if the latter is conceptually expanded from meaning system imperfections to also covering imperfect user memory.
@@ 5.3.2. Memory vs Text Cues
5.4. Limitations, Generalisability, and Future work
Here, we discuss limitations, generalisability of our findings, and future research directions.
5.4.1. Writing Tasks
We only examined the impact of AI usage during ideation and elaboration tasks as part of the writing process. We opted for those tasks as they represent fundamental stages that precede and influence many other creative processes. However, many real-world uses of LLMs involve different cognitive demands and output structures. Future work could therefore examine whether similar AI-induced source confusion occurs in tasks such as summarisation, paraphrasing, document search, information extraction, or multimodal analysis.
We hypothesise that these tasks differ in ways that may widen or close the AI memory gap. For instance, summarisation and paraphrasing often build directly on previous text and may blur the boundary between self- and AI-generated content even further, especially in mixed workflows. Remembering whether one conducted document or information search may also be more difficult: AI-mediated search removes users from the step-by-step exploration that normally creates distinctive memory cues (e.g., navigating papers or websites), and AI systems frequently pre-extract relevant information directly in the chat window, further reducing contextual markers. Conversely, tasks that require substantial transformation or personal reasoning may preserve stronger cues of self-authorship.
Related, whether our findings apply to domains other than writing, such as coding, image and video creation, or design, needs to be explored in future work. In addition, our experimental task mimicked a co-creative workflow for all AI conditions, where the human always made the final decision. Situations in which AI makes executive decisions on its own remain open for investigation. Based on our results, we hypothesise that this would further reduce people’s ability to remember parts of the creative process. A needed next step is to identify which task characteristics and interface cues most strongly shape the AI memory gap, with a study in the field and with full writing workflows.
5.4.2. AI Interactions
Our study focused on a small subset of possible AI interactions, specifically using an LLM (GPT-4.1-mini) through a chatbot interface. Using different models, interaction paradigms (e.g., inline suggestions), or interfaces could lead to different outcomes. Moreover, participants completed the Phase 2 memory tasks without access to any external aids, such as prompt histories or chat logs. This was intentional, as our aim was to measure source memory ability in isolation. However, in practice, users may have access to persistent chat histories, document revision trails, or retrieval features. Such aids do not necessarily improve memory directly, but they do preserve interaction traces that users can revisit, which may change how source attribution is solved in a specific practical context.
5.4.3. Ecological Factors
In our study, people created ideas in one session, whereas typical creative workflows often stretch over multiple sessions. We also tested people’s source memory after a fixed time gap of approximately one week, and it remains unclear how shorter or longer retention intervals might influence memory performance in this context. Moreover, individuals’ motivation and effort to recall AI contributions may vary in different real-world settings, for example when people feel pressure to justify their choices in a group setting or when they are simply exploring ideas informally for personal inspiration. Future work could also examine memory performance when participants are informed in advance that they will later be asked to report their use of AI assistance.
@@ A.1. Problem Statements
The known-topic and unknown-topic distractors can be found in our project repository (https://osf.io/dvkj9/).